# Can Synthetic Images Serve as Effective and Efficient Class Prototypes?

### v1

@@ Abstract
Vision-Language Models (VLMs) have shown strong performance in zero-shot image classification tasks. However, existing methods, including Contrastive Language-Image Pre-training (CLIP), all rely on annotated text-to-image pairs for aligning visual and textual modalities. This dependency introduces substantial cost and accuracy requirement in preparing high-quality datasets. At the same time, processing data from two modes also requires dual-tower encoders for most models, which also hinders their lightweight. To address these limitations, we introduce a “Contrastive Language-Image Pre-training via Large-Language-Model-based Generation (LGCLIP)” framework. LGCLIP leverages a Large Language Model (LLM) to generate class-specific prompts that guide a diffusion model in synthesizing reference images. Afterwards these generated images serve as visual prototypes, and the visual features of real images are extracted and compared with the visual features of these prototypes to achieve comparative prediction. By optimizing prompt generation through the LLM and employing only a visual encoder, LGCLIP remains lightweight and efficient. Crucially, our framework requires only class labels as input during whole experimental procedure, eliminating the need for manually annotated image-text pairs and extra pre-processing. Experimental results validate the feasibility and efficiency of LGCLIP, demonstrating great performance in zero-shot classification tasks and establishing a novel paradigm for classification.
@@ 1 Introduction
In recent years, deep learning has profoundly transformed computer vision, with the emergence of Vision-Language Models (VLMs) marking a significant paradigm shift [guo2023calip]. Unlike traditional supervised learning methods that rely heavily on labeled data, VLMs learn from large-scale image-text pairs and exhibit unprecedented generalization abilities [udandarao2023sus]. Among them, CLIP [radford2021learning] has achieved remarkable success by aligning image and text features in a shared embedding space, enabling natural language supervision for visual tasks. By simply converting category labels into textual prompts (e.g., “a photo of a class”), CLIP achieves zero-shot image classification (ZSL) without requiring task-specific training data, thereby reducing dependence on labeled datasets [chen2025interpretable, qian2024online].
Despite these advances, the core paradigm of image-text alignment presents inherent limitations. Performance depends heavily on the wording of textual prompts, which has led to costly and unstable “prompt engineering” [saha2024improved, zhang2022tip]. Automated solutions such as CoOp[Zhou_2022] and CoCoOp [Zhou_2022_CVPR] learn soft prompts to adapt across tasks, but they suffer from overfitting, reduced generalization to novel categories, and high computational cost due to optimization through large encoders. Consequently, reliance on text encoders during inference remains a fundamental bottleneck [zhang2022show].
In parallel, text-to-image (T2I) diffusion models have advanced rapidly, demonstrating the ability to generate highly realistic and diverse images [he2023syntheticdatagenerativemodels]. These models have been applied successfully to augment training datasets with synthetic samples, improving classifier robustness and zero-shot ability [azizi2023syntheticdatadiffusionmodels]. Recent research further combines diffusion models with large language models (LLMs) to overcome the limitations of simple category prompts. LLMs can generate detailed and unbiased descriptions [qu2023layoutllm], guiding diffusion models to synthesize higher-quality and more controllable images. Such LLM-guided generation has been shown to improve data diversity and reduce semantic bias, thereby enhancing downstream classification performance [qin2024diffusiongpt].
@@ 2 RELATED WORK
VLM Architectures. CogVLM[CogVLM] integrates a trainable ”visual expert module” at each Transformer layer to solve the problem of insufficient fusion between visual and language features in shallow alignment architectures. LLaVA[LLaVA] uses a simple linear layer to project visual features into the language model’s word embedding space, effectively connecting pre-trained vision and language models with minimal training overhead for visual instruction fine-tuning.
Optimizing Text-Image pairs for images. CoDA[CoDA] uses contrastive visual data augmentation to guide the diffusion model in generating targeted synthetic images, addressing poor performance in recognizing novel or confused concepts due to data insufficiency . MosaicFusion[mosaicfusion] uses a diffusion model to generate multiple controllable objects and segmentation masks across different regions of a single image canvas, solving the issue of limited labeled data in instance segmentation tasks, especially in long-tail and open-set scenarios .
@@ 3.2 Diffusion model generates images
After obtaining the processed prompts, we employ a diffusion model to generate images. Among various generative approaches—such as GANs, VAEs, and diffusion models—we adopt Stable Diffusion (SD) for its balanced trade-off between controllability, efficiency, and image quality. Consistent with our core motivation of “using unbiased generated images as the basis for classifying real images,” we select SD 2.1 as the generation engine, as it produces realistic-style images with strong text-to-image capability [onwurah2025knowing]. To ensure reproducibility, a fixed global random seed is applied.
@@ 4.1 Dataset
To conduct a comprehensive evaluation, we selected six widely-used datasets for zero-shot learning, encompassing a broad range of objects, scenes, and fine-grained categories: CUB [CUB], FLO [FLO], PET [PET], ImageNet [imagenet], FOOD [FOOD], and EUROSAT [eurosat]. These datasets serve as the foundational data for the entire experimental process.Since the training process is not involved, we use all the original images from each dataset as the test set.
@@ 4.2 Experimental Setup
Feature extraction and contrastive prediction. Visual features are extracted using CLIP backbones (RN50 [RN50], ViT-B/32, ViT-B/16, ViT-L/14 [ViT]). For an image $I_{p}$, we compute a multi-scale weighted feature $F_{s}^{p}$. Features from $N_{g}$ images are sampled and average-pooled to form the prototype of class $c$. A contrastive score matrix assigns each real image to the class with highest similarity, and final performance is reported using macro accuracy.
@@ 4.5 Limitations and expectations
There is still a gap in accuracy compared to traditional methods like CLIP’s T2I method [radford2021learning], and the performance varies greatly across different datasets. This is the main deficiency. We look forward to optimizing each aspect of our method in the future to achieve performance comparable to traditional methods.

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown strong performance in zero-shot image classification tasks. However, existing methods, including Contrastive Language-Image Pre-training (CLIP), all rely on annotated text-to-image pairs for aligning visual and textual modalities. This dependency introduces substantial cost and accuracy requirement in preparing high-quality datasets. At the same time, processing data from two modes also requires dual-tower encoders for most models, which also hinders their lightweight. To address these limitations, we introduce a “Contrastive Language-Image Pre-training via Large-Language-Model-based Generation (LGCLIP)” framework. LGCLIP leverages a Large Language Model (LLM) to generate class-specific prompts that guide a diffusion model in synthesizing reference images. Afterwards these generated images serve as visual prototypes, and the visual features of real images are extracted and compared with the visual features of these prototypes to achieve comparative prediction. By optimizing prompt generation through the LLM and employing only a visual encoder, LGCLIP remains lightweight and efficient. Crucially, our framework requires only class labels as input during whole experimental procedure, eliminating the need for manually annotated image-text pairs and extra pre-processing. Experimental results validate the feasibility and efficiency of LGCLIP, demonstrating great performance in zero-shot classification tasks and establishing a novel paradigm for classification. Codes at Project Link.
00footnotetext: Project lead; Corresponding author.
@@ 1 Introduction
In recent years, deep learning has profoundly transformed computer vision, with the emergence of Vision-Language Models (VLMs) marking a significant paradigm shift [8]. Unlike traditional supervised learning methods that rely heavily on labeled data, VLMs learn from large-scale image-text pairs and exhibit unprecedented generalization abilities [19]. Among them, CLIP [17] has achieved remarkable success by aligning image and text features in a shared embedding space, enabling natural language supervision for visual tasks. By simply converting category labels into textual prompts (e.g., “a photo of a class”), CLIP achieves zero-shot image classification (ZSL) without requiring task-specific training data, thereby reducing dependence on labeled datasets [4, 14].
Despite these advances, the core paradigm of image-text alignment presents inherent limitations. Performance depends heavily on the wording of textual prompts, which has led to costly and unstable “prompt engineering” [18]. Automated solutions such as CoOp[25] and CoCoOp [24] learn soft prompts to adapt across tasks, but they suffer from overfitting, reduced generalization to novel categories, and high computational cost due to optimization through large encoders. Consequently, reliance on text encoders during inference remains a fundamental bottleneck [23].
In parallel, text-to-image (T2I) diffusion models have advanced rapidly, demonstrating the ability to generate highly realistic and diverse images [10, 7]. These models have been applied successfully to augment training datasets with synthetic samples, improving classifier robustness and zero-shot ability [1]. Recent research further combines diffusion models with large language models (LLMs) to overcome the limitations of simple category prompts. LLMs can generate detailed and unbiased descriptions [16], guiding diffusion models to synthesize higher-quality and more controllable images. Such LLM-guided generation has been shown to improve data diversity and reduce semantic bias, thereby enhancing downstream classification performance [15].
@@ 2 RELATED WORK
VLM Architectures. CogVLM[21] integrates a trainable ”visual expert module” at each Transformer layer to solve the problem of insufficient fusion between visual and language features in shallow alignment architectures. LLaVA[13] uses a simple linear layer to project visual features into the language model’s word embedding space, effectively connecting pre-trained vision and language models with minimal training overhead for visual instruction fine-tuning.
Optimizing Text-Image pairs for images. CoDA[26] uses contrastive visual data augmentation to guide the diffusion model in generating targeted synthetic images, addressing poor performance in recognizing novel or confused concepts due to data insufficiency . MosaicFusion[22] uses a diffusion model to generate multiple controllable objects and segmentation masks across different regions of a single image canvas, solving the issue of limited labeled data in instance segmentation tasks, especially in long-tail and open-set scenarios .
@@ 3.2 Diffusion model generates images
After obtaining the processed prompts, we employ a diffusion model to generate images. Among various generative approaches—such as GANs, VAEs, and diffusion models—we adopt Stable Diffusion (SD) for its balanced trade-off between controllability, efficiency, and image quality. Consistent with our core motivation of “using unbiased generated images as the basis for classifying real images,” we select SD 2.1 as the generation engine, as it produces realistic-style images with strong text-to-image capability. To ensure reproducibility, a fixed global random seed is applied.
@@ 4.1 Dataset
To conduct a comprehensive evaluation, we selected six widely-used datasets for zero-shot learning, encompassing a broad range of objects, scenes, and fine-grained categories: CUB [20], FLO [3], PET [12], ImageNet [5], FOOD [2], and EUROSAT [11]. These datasets serve as the foundational data for the entire experimental process.Since the training process is not involved, we use all the original images from each dataset as the test set.
@@ 4.2 Experimental Setup
Feature extraction and contrastive prediction. Visual features are extracted using CLIP backbones (RN50 [9], ViT-B/32, ViT-B/16, ViT-L/14 [6]). For an image $I_{p}$, we compute a multi-scale weighted feature $F_{s}^{p}$. Features from $N_{g}$ images are sampled and average-pooled to form the prototype of class $c$. A contrastive score matrix assigns each real image to the class with highest similarity, and final performance is reported using macro accuracy.
@@ 4.5 Limitations and expectations
There is still a gap in accuracy compared to traditional methods like CLIP’s T2I method [17], and the performance varies greatly across different datasets. This is the main deficiency. We look forward to optimizing each aspect of our method in the future to achieve performance comparable to traditional methods.